please help me solve this equation: 4/y+5y=21/2

Answer by mananth(16946)

4/y +5y =21/2
4+5y^2 / y = 21/2
2(4+5y^2)=21y
8+10y^2=21y
10y^2-21y+8=0
10y^2-5y-16y +8=0
5y(2y-1)-8(2y-1)=0
(2y-1)(5y-8)=0
y=1/2 OR y= 8/5
Ananth